jim is 5 years older than cindy.4 years ago, the sum of their ages was 21.find each of their ages now
Answer by simranb(55)

Let cindy's age be x
Then jim's age= x+5
Cindy's age 4 years ago=x-4
Then jim's age 4 years ago = x+5-4
= x+1
Sum of their ages 4 years ago =21
Equation
x-4+x+1=21
2x-3=21
2x-3=21
2x=21+3
2x=24
x=24/2
x=12
Cindy's age=12yrs
Jim's Age = 12+5
= 17yrs
Cheers!
Hope this helps!